OpenStudy (anonymous):

f(n + 1) = f(n) – 3. If f(4) = 22, what is f(1)?

OpenStudy (anonymous):

f(n + 1) is the term after f(n). You're subtracting 3 to get to the next term. So work backwards and add 3 to get the previous term. f(4) = 22 f(3) = 22 + 3 = 25 f(2) = 25 + 3 and so on